The average age of 6 men is 49 years and the average of 5 of them is 47 years.what is the average of the sixth man?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52782)

,

The total age of the 6 men is  6*49 = 294 years.


The total age of the 5 of them is  5*47 = 235 years.


The difference,  294 - 235 = 59 years is the age of the sixth man.    ANSWER

Solved.

-------------------

To see many other similar solved problems,  look into the lessons
    - Solved problems on average scores, weight, height and temperature
    - Solved problems on average scores
    - Solved problems on average age
    - Miscellaneous problems on average values
in this site.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.


Happy New Year (!)



Answer by greenestamps(13200)



Here is a way to solve this kind of problem that is different than the usual method shown by the other tutor. It can be useful when the numbers being averaged are all close together.

The average age of all 6 men is 49, and the average age of 5 of them is 47.

Comparing the known ages to the average, the age of each of the 5 men is 2 years below the average, so the total age of those 5 men is 5*2=10 years under the average.

That means the age of the 6th man must be 10 years above the average.

ANSWER: The age of the 6th man is 49+10 = 59.